# Attention In Geometry: An Adaptive Density Field (ADF) with FAISS-Accelerated Spatial Attention

### v1

@@ 2.6.3 Why This Matters
While ADF shares mathematical components with adaptive KDE and ANN-accelerated density estimation, its contribution is not a new kernel estimator per se. Instead, ADF explicitly formulates these components as a geometric attention operator, where approximation, sparsification, and metric structure are intrinsic to the definition rather than implementation details.
Importantly, the attention analogy is used as a structural lens rather than a claim of architectural novelty. ADF does not seek to replace or subsume Transformer-style attention, but to extend the abstract notion of query-conditioned weighted aggregation to continuous metric spaces.
Overall, ADF should be understood as a framework-level operator rather than a fixed algorithm, with emphasis on structure, scalability, and geometric grounding.
3. Implementation and results
Given a large set of aircraft trajectories, local motion irregularities are first detected as Points of interest (POIs). These POIs are sparse, noisy, and heterogeneous. We aim to aggregate them into a continuous spatial field to reveal Zones of Influence (ZOIs): regions where systematic maneuvering, airspace structure, or control-induced behavior recurrently concentrates.
In many spatial decision problems (e.g., flight routing, navigation, risk exposure), the agent does not respond to absolute field intensity, but to relative dominance along its trajectory. Zones of Influence (ZOIs) are therefore introduced as trajectory-conditioned dominance regions, representing spatial areas where the accumulated influence field locally dominates competing influences relative to the agent’s path.
3.1. Study area and data description
We used a nationwide flight trajectory dataset over mainland China for two multi-periods: 2024-11-10 to 2024-11-22 and 2024-12-05 to 2024-12-15 to extract the POIs. There are a total of 1.8 millions of POIs extracted from the dataset using the motion-based procedure described in Section 3.2.
For the case study on trajectory-conditioned ZOIs, we selected a single day and airport: the Chengdu Shuangliu airport (CTU) on 2024-12-16. All trajectories intersecting the Chengdu region on that day were used to construct a local example of the Adaptive Density Field (ADF) and to illustrate ZOI extraction along individual flights.
The ADF framework is not restricted to flight trajectories and can be instantiated in any domain where localized motion irregularities or events are observed. Likewise, the POI definition is not universal; the construction adopted here is a concrete example tailored to aircraft motion.
3.2. POI construction
To collect the POI data, we applied a mixture of motion-based POI detection methods, detailed in the Appendix. However, POIs are derived from trajectory segments where a physics‑based motion model fails to predict future positions accurately. The processing pipeline operates at the individual flight level. For each flight, raw WGS‑84 positions (longitude, latitude, altitude) are first converted to a local East–North–Up (ENU) coordinate system. The trajectory is then represented as a dictionary keyed by flight identifier, with three aligned time series:
  • coords: longitude, latitude, altitude
  • vel: velocity components (velocity on each axis)
  • dt: timestamp interval
To estimate future aircraft positions, I applied a physics-based interpolation model that blends two motion predictors:
@@ 3.2. POI construction
the blending weight is determined by the local curvature of the trajectory [lee2007trajectory]: Low curvature indicates motion is nearly straight, hence CA weights more, and vice versa. This adaptive combination produces a more stable and realistic prediction thanm using either method alone.
To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss (This metric captures both the magnitude and directional structure of prediction errors.) for each flight. This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing. After computing the time‑normalized Mahalanobis loss for each flight, the next step is to identify Points of Interest (POIs)—locations where the prediction error is unusually high [chandola2009anomaly]. These points often correspond to sharp maneuvers, abnormal motion, or sensor irregularities, and they serve as valuable markers for downstream analysis.
However, in this case, a POI does not necessarily correspond to an actual infrastructure feature; it simply marks a point where the motion deviates significantly from our prediction. It can be regard as an example tailored POI definition under certain circumstances.
Later we normalize the loss scores and filter the POI with a threshold and exported the data into csv format.
3.3. ADF construction
We instantiated the ADF framework on the nationwide POI dataset by specifying the coordinate system, kernel parameterizations, and neighbor search configurations. All POI locations were converted from WGS84 geodetic coordinates to Earth-Centered, Earth-Fixed (ECEF) Cartesian coordinates to ensure metric consistency. POI scores were treated as non-negative scalar probability attention values and were used directly without additionaly normalization.
For the scalable neighbor retrieval, a FAISS inverted file (IVF) index was introduced and constructed over the ECEF coordinates. The index was trained with 4096 clusters ($2^{12}$) clusters, and queries probed the 16 nearest clusters. For each query point the top $k=100$ approximate nearest neighbors were retrieved. The global spatial scale parameter was fixed to $\sigma_{0}=500$ meters for all experiments. ADF values were evaluated using the formulation defined in Section 2, without further modification.
3.4. Field evaluation
To analyze the ADF along individual trajectories in Chengdu, we:
@@ 3.4. Field evaluation
    Convert the coordinates: transform all points from geodetic to ECEF.
@@
    Evaluate ADF: compute $ADF(x_{t})$ at each trajectory point $x_{t}$ using the FAISS-accelerated framework.
  3. 3.
    Re-attach geodetic coordinates: join ADF values back to $(lon,lat,alt)$ for downstream analysis.
This allows us to examine how aircraft trajectories intersect regions of concentrated POIs.
3.5. ZOI extraction
Given the ADF evaluated along a trajectory, we extract Zones of Influence (ZOIs) as segments where the field intensity dominates a baseline [zheng2015trajectory] defined by the flight’s own experience. Let $\{F_{t}\}$ denote the sequence of ADF values along a single trajectory. We define a robust baseline level as the trajectory‑wise median [rousseeuw1993alternatives]:
$F_{base}={median}_{t}(F_{t})$
A point $t$ is labeled as belonging to a ZOI if
$F_{t}\geq\alpha F_{base}$
where $\alpha>1$ is a dominance factor. In the Chengdu case study, we set $\alpha=1.3$, chosen empirically to balance sensitivity and specificity in detecting high-density segments along individual trajectories. The resulting binary mask partitions the trajectory into ZOI and non‑ZOI segments. Trajectories primarily in low-density regions may still exhibit short ZOIs, while trajectories in busy airspace require higher ADF values to exceed the baseline.
Figure 2: Trajectory ADF analysis and ZOI extraction
Visualization
To illustrate the ADF and ZOI extraction:
  • Trajectories are rendered as white lines, tracing the actual flight paths through the airspace.
  • Red dots mark the extracted Zones of Influence (ZOIs)—trajectory points where the ADF value exceeds the dominance threshold relative to the flight’s own baseline.
  • The remaining trajectory points are colored by their Adaptive Density Field (ADF) values (within range 0~30), using a continuous colormap from blue (low intensity) to red (high intensity), with values ranging from 0 to 80.
  • Layered maps allow toggling between POIs, ADF, ZOIs, and trajectory paths, enabling flexible inspection of each component.
These visualizations demonstrate the relationship between the continuous ADF field and trajectory-conditioned ZOIs. High-ADF segments generally coincide with operationally significant areas such as approach corridors, holding patterns, and vectoring zones in the Chengdu terminal. ZOI points, highlighted in red, mark locations where the trajectory locally exceeds its baseline field intensity; some of these points overlap dense ADF regions, representing recurrent maneuvering, while others appear in otherwise sparse areas, potentially reflecting transient anomalies, unique pilot actions, or data artifacts. By layering ADF intensity, ZOI labels, and raw POIs, the visualization distinguishes between structurally recurrent behavior and isolated events, offering a nuanced perspective on spatial influence and motion irregularities.
4. Discussion
The proposed Adaptive Density Field (ADF) framework provides a flexible, scalable mechanism for aggregating sparse, heterogeneous points of interest (POIs) into a continuous spatial field. By coupling trajectory-conditioned analysis with ZOI extraction as a downstream instantiation of ADF, this methodology captures both globally recurrent patterns and locally significant deviations.
4.1. Future application and extensions
  1. 1.
    Operational Airspace Management: The ADF-ZOI pipeline we presented in the example implementation section could assist air traffic controllers in identifying target regions depending on the specific POI definition, such as regions of frequent maneuvering or congestion, enabling data-driven decisions for route planning, holding pattern optimization, and as well as safety monitoring.
  2. 2.
    Predictive Trajectory Analysis: Beyond post-hoc evaluation, the framework can be integrated with predictive models for trajectory planning, anomaly detection, or risk exposure assessment. By defining the POIs, it is possible to derive trajectory-conditioned ZOIs to provide adaptive thresholds for identifying unusual events relative to expected motion patterns.
  3. 3.
    Cross-Domain Applicability: While this study focuses on aircraft trajectories due to the initial problem setting, the methodolody is domain-agnostic. Any scenario involving spatiotemporal events, as long as the spatial characteristics are stable and fixed (in other words, not dynamic), such as maritime traffic, pedestrian movement, vehicle flows, or wildlife tracking – we expect ADF-based spatial attention is able to identify zones of concentrated activities.
  4. 4.
    Integration with Semantic Context: Future work could embed semantic labels, such as airspace type, weather conditions, or operational procedures, into the ADF formulation, yielding geo-semantic density fields. This way, interestingly, aligns even more closely with the application of attention mechanism in spatiotemporal analysis.
  5. 5.
    Real-Time Applications: With optimizations to indexing, neighbor search, or GPU-accelerated computation, the ADF framework could support near-real-time monitoring of live trajectories in drones, robotics, and other autonomous systems, allowing dynamic ZOI identification and interactive visualization for operational decision-making.
4.2. Methodological Reflections
  • Parameter choices, such as kernel bandwidth and dominance factor, influence sensitivity and specificity; future studies could explore adaptive or data-driven tuning to optimize ZOI detection across heterogeneous datasets.
  • Visualization remains crucial for interpretation: layered maps combining POIs, ADF intensity, and ZOI labels provide immediate insight into both recurrent behavior and outlier events, but higher-dimensional visualizations or interactive dashboards could further enhance understanding.
5. Conclusion
This work introduces Adaptive Density Fields (ADF) as a formulation-level geometric attention operator for scalable aggregation of sparse spatial events in continuous metric space. By defining spatial influence through query-conditioned neighbor selection, score-modulated kernel weighting, and approximate nearest-neighbor execution, ADF reframes adaptive kernel aggregation as an intrinsic form of geometry-preserving attention rather than a purely statistical estimator. The primary contribution lies not in a specific kernel choice or application, but in the structural decomposition of spatial aggregation: locality induced by k-NN attention, metric-grounded weighting, and approximation treated as a defining component of the operator. This perspective unifies ideas from adaptive kernel methods, spatial indexing, and attention mechanisms under a common geometric framework that scales to millions of points while remaining interpretable. To demonstrate the utility of this formulation, we presented an example instantiation in the context of aircraft trajectory analysis. Motion-derived Points of Interest were aggregated into a continuous ADF, and trajectory-conditioned Zones of Influence (ZOIs) were extracted using a relative dominance criterion along individual paths. The Chengdu case study illustrates how the resulting fields reveal recurrent airspace structures while distinguishing localized, trajectory-specific deviations. Beyond aviation, the ADF framework is domain-agnostic and applicable to a wide range of spatial and spatiotemporal settings, including urban mobility, maritime traffic, robotics, and environmental monitoring. Future work will explore semantic augmentation of the field, adaptive parameter learning, anisotropic kernels, and real-time deployment, further extending the applicability of geometric attention in continuous spatial systems. Overall, ADF provides a robust and extensible foundation for scalable, interpretable spatial attention, bridging geometric structure and efficient computation in complex spatial environments.
@@ 5. Conclusion
1 Preprocessing the Flight Data
1.1 Coordinate Conversion [hofmann2012global]: WGS84 Geodetic to ECEF
@@
1.2 Coordinate Conversion: ECEF to ENU Conversion
@@ 2 Position Prediction
    Constant‑Acceleration (CA) model [bar2001estimation] — reliable for nearly straight trajectories
@@
    Cubic Hermite Spline interpolation [de1978practical] — smooth and accurate for curved motion
The spline utilizes local velocity vectors as tangents at each waypoint, providing a geometrically consistent path [virtanen2020scipy] that complements the CA model’s acceleration-based predictions."
@@ 2.1 General Prediction
2.2 Loss Computation
To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss for each flight. This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing.
@@ 2.2 Loss Computation
1. Extract Prediction Residuals
@@
2. Estimate Residual Covariance
@@
A small Tikhonov regularization [tikhonov1963solution] term $\lambda=10^{-5}$ stabilizes the inversion of $\Sigma$, especially for nearly collinear trajectories.
@@
3. Compute Mahalanobis Distance
@@
4. Normalize by Temporal Spacing
@@ 3 POI Detection
3.1 Normalize the Loss Scores

### v2

@@ 5.5 Why This Matters
The suggested ADF is best understood as a geometric attention mechanism: a query‑conditioned, metric‑driven aggregation over continuous space. In this sense, ADF implements a spatial attention operator rather than a traditional kernel estimator. While ADF shares mathematical components with adaptive KDE and ANN-accelerated density estimation, its contribution is not a new kernel estimator per se. Instead, ADF explicitly formulates these components as a geometric attention operator, where approximation, sparsification, and metric structure are intrinsic to the definition rather than implementation details. Overall, ADF should be understood as a framework-level operator rather than a fixed algorithm, with emphasis on structure, scalability, and geometric grounding.
6 Discussion
The proposed Adaptive Density Field (ADF) framework provides a flexible, scalable mechanism for aggregating sparse, heterogeneous points of interest (POIs) into a continuous spatial field. By coupling trajectory-conditioned analysis with POI extraction as a downstream instantiation of ADF, this methodology captures both globally recurrent patterns and locally significant deviations.
6.1 Future application and extensions
@@
    Operational Airspace Management: The ADF-POI pipeline we presented in the Section 3 could assist air traffic controllers in identifying target regions depending on the specific POI definition, such as regions of frequent maneuvering or congestion, enabling data-driven decisions for route planning, holding pattern optimization, and and safety monitoring.
@@
    Predictive Trajectory Analysis: Beyond post-hoc evaluation, the framework can be integrated with predictive models for trajectory planning, anomaly detection, or risk exposure assessment. By defining the POIs, it is possible to derive trajectory-conditioned POIs to provide adaptive thresholds for identifying unusual events relative to expected motion patterns.
  3. 3.
    Cross-Domain Applicability: While this study focuses on aircraft trajectories due to the initial problem setting, the methodology is domain-agnostic. Any scenario involving discrete spatiotemporal observations with underlying metric regularities, as long as the spatial characteristics are stable and fixed (in other words, not dynamic), such as maritime traffic, pedestrian movement, vehicle flows, or wildlife tracking – we expect ADF-based spatial attention can identify zones of concentrated activities.
  4. 4.
    Integration with Semantic Context: Future work could embed semantic labels, such as airspace type, weather conditions, or operational procedures, into the ADF formulation, yielding geo-semantic density fields. This way, interestingly, aligns even more closely with the application of the attention mechanism in spatiotemporal analysis.
  5. 5.
    Real-Time Applications: With optimizations to indexing, neighbor search, or GPU-accelerated computation, the ADF framework could support near-real-time monitoring of live trajectories in drones, robotics, and other autonomous systems, allowing dynamic POI identification and interactive visualization for operational decision-making.
6.2 Methodological Reflections
  • Parameter choices, such as kernel bandwidth and dominance factor, influence sensitivity and specificity; future studies could explore adaptive or data-driven tuning to optimize POI detection across heterogeneous datasets.
  • Visualization remains crucial for interpretation: layered maps combining POIs, ADF intensity, and POI labels provide immediate insight into both recurrent behavior and outlier events, but higher-dimensional visualizations or interactive dashboards could further enhance understanding.
7 Conclusion
This work introduces Adaptive Density Fields (ADF) as a formulation-level geometric attention operator for scalable aggregation of sparse spatial events in continuous metric space. By defining spatial influence through query-conditioned neighbor selection, score-modulated kernel weighting, and approximate nearest-neighbor execution, ADF reframes adaptive kernel aggregation as an intrinsic form of geometry-preserving attention rather than a purely statistical estimator. The primary contribution lies not in a specific kernel choice or application, but in the structural decomposition of spatial aggregation: locality induced by k-NN attention, metric-grounded weighting, and approximation treated as a defining component of the operator. This perspective unifies ideas from adaptive kernel methods, spatial indexing, and attention mechanisms under a common geometric framework that scales to millions of points while remaining interpretable.
To demonstrate the utility of this formulation, we presented an example instantiation in the context of aircraft trajectory analysis. Motion-derived Points of Interest were aggregated into a continuous ADF, and trajectory-conditioned Points of Interest (POIs) were extracted using a relative dominance criterion along individual paths. The case study illustrates how the resulting fields reveal recurrent airspace structures while distinguishing localized, trajectory-specific deviations. Beyond aviation, the ADF framework is domain-agnostic and applicable to a wide range of spatial and spatiotemporal settings, including urban mobility, maritime traffic, robotics, and environmental monitoring. Future work will explore semantic augmentation of the field, adaptive parameter learning, anisotropic kernels, and real-time deployment, further extending the applicability of geometric attention in continuous spatial systems. Overall, ADF provides a robust and extensible foundation for scalable, interpretable spatial attention, bridging geometric structure and efficient computation in complex spatial environments.
Notes
@@ 7 Conclusion
    The raw dataset used in the experiments is proprietary and cannot be made public.
@@
    All code is publicly available at: https://github.com/JaimeFine/adaptive-density-field.
Acknowledgements
The author would like to thank Dr. Yunxiang Han (Sichuan University) for valuable guidance and constructive feedback during the development of this work. This research also made use of open-source scientific software, including Python and R for data processing and analysis, and the Leaflet library with OpenStreetMap basemaps (© OpenStreetMap contributors) for interactive geospatial visualization. Figures and manuscript preparation were conducted using .
Generative artificial intelligence
This manuscript benefited from generative AI tools in limited, non-substantive ways. ChatGPT (version 4o-mini) and Gemini 3 were used to improve language clarity and fluency. Microsoft Copilot (search mode) was employed to assist with citation discovery and verification. All conceptual content, analysis, and argumentation were developed by the author.
@@ Acknowledgements
1. Preprocessing the Flight Data
1.1. Coordinate Conversion [16]: WGS84 Geodetic to ECEF
@@
1.2. Coordinate Conversion: ECEF to ENU Conversion
@@ 2. Position Prediction
    Constant‑Acceleration (CA) model [3] — reliable for nearly straight trajectories
@@
    Cubic Hermite Spline interpolation [11] — smooth and accurate for curved motion
The spline utilizes local velocity vectors as tangents at each waypoint, providing a geometrically consistent path [36] that complements the CA model’s acceleration-based predictions."
@@ 2.1. General Prediction
2.2. Loss Computation
To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss [22] for each flight. This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing.
@@ 2.2. Loss Computation
Extract Prediction Residuals
@@
Estimate Residual Covariance
@@
A small Tikhonov regularization [34] term $\lambda=10^{-5}$ stabilizes the inversion of $\Sigma$, especially for nearly collinear trajectories.
@@
Compute Mahalanobis Distance
@@
Normalize by Temporal Spacing
@@ 3. POI Detection
3.1. Normalize the Loss Scores